OpenStudy (anonymous):

7+sqrt (x-7)=9 Please help me solve this

OpenStudy (doc.brown):

? \[7+\sqrt{x-7}=9\]

OpenStudy (doc.brown):

Square everything

OpenStudy (anonymous):

Isolate x. Subtract 7 from both sides: \[7-7+\sqrt{x-7}=9-7\] \[\sqrt{x-7} = 2\] Square both sides: \[(\sqrt(x-7))^2=2^2\] \[(x-7)=4\] Add 7 to both sides: \[x-7+7 = 4 + 7\] Simplify: \[x = 11\]

OpenStudy (doc.brown):

Yeah, like that.